OpenStudy (anonymous):

what is y - 1 = -4/5(x-3) written in standard form using positive integers

OpenStudy (anonymous):

and i don't know how.......

OpenStudy (anonymous):

well you can start with the fraction side - ( 4/5 ) (x-3)

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

Ok i did....and i get 4/5x + y = 3.4

OpenStudy (anonymous):

4/5 = 0.8

OpenStudy (anonymous):

Do do I round that to 1? that is one of my answers.

OpenStudy (anonymous):

emm well no

OpenStudy (anonymous):

graphing you dont round things off

OpenStudy (anonymous):

well, i'm not graphing this

OpenStudy (anonymous):

i mean this is slope intercept form meaning to give directions to graph

OpenStudy (anonymous):

hm...well that's a problem

OpenStudy (anonymous):

well its easier for you you dont have to round things off so you can leave it like that providing it is correct

OpenStudy (anonymous):

nope....0.8 is not one of the answers

OpenStudy (anonymous):

what you have right there is the point slope form what hat you can get out of it is the slope intercept formula <----- the answer you want

OpenStudy (anonymous):

you want a formula for your answer

OpenStudy (anonymous):

a formula? i only have numbers for answers 4 5 3 1

OpenStudy (anonymous):

it should be in the form y = mx + b y = y value x = x value m = slope b = y intercept where the line intersects y

OpenStudy (anonymous):

ok one sec lemme figure that out

OpenStudy (anonymous):

and i get y = -4/5x + 12/5 (or 2.4)

OpenStudy (anonymous):

or, do i plug in the numbers i have for an answer and see if it comes out correct?

OpenStudy (anonymous):

However, I want it in standard form, which is Ax + Bx = C

OpenStudy (anonymous):

gtg sorry sleep :/

OpenStudy (anonymous):

oh...ok

OpenStudy (anonymous):

oh ya about just switch it around/ my mistake

OpenStudy (anonymous):



OpenStudy (anonymous):

algebraically

OpenStudy (anonymous):

im off gn :)

OpenStudy (mandre):

The only 2 standard forms I know of are y = mx + b and Ax + By = C In either form you get y = -4/5x + 17/5 or 4x + 5y = 17 (simplified to integer form). I don't see where the answer can be an integer. Strange question.